SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (April 4, 2006)

Attendees: Rainer Bartoldus, Kris Chaisanguanthum, Al Eisner, Fang Fang, Selina Li, Rahmat Rahmat, Dong Su

Operations

  • We are taking physics data again after the shutdown for Q2 chamber/bellows replacement. PEP was able to raise the LER current high beyond before the shutdown and no vacuum burst observed. It looks like the vacuum burst problem was resolved.
  • The TSF-X crate seems to cause some problems. See the DCT/DCZ section below.

EMT

  • The EMT deadtime is about 0.8% on average. It looks like the FEX suppression didn't help us as much as anticipated, and perhaps there is something else such as the fiber transport that causes it. We will have a discussion with Chris O'Grady on Friday after the detector operation meeting to see what we can do about the deadtime.

DCT/DCZ

  • On March 29 (Wednesday) and April 3 (Monday), we had to power cycle the TSF-X crate when DCZ failed to configure. This might be associated with the DCM due to the new TSF firmware. We wonder if the delay in the configure is set properly or long enough. Rainer suggested maybe to put the delay into user reset.
  • There were out-of-order and out-of-synch error messages in the TSF-X crate on April 3 (for ~4 seconds) and April 4 (for ~1 minute and 7 seconds). Shipping out the input datagram would help diagnostics. Rainer will help Fang with the online code change.
  • We discuss turning off the old DCT system. Chris O'Grady gave us the confirmation that DCH/ODF no longer need the old DCT as far as they concern. We will leave the old DCT system out of data taking, but still in such a state that we could turn it back on if necessary for some amount of time till the DCM issue mentioned above is completely resolved.

DCZ online

  • Fang is cleaning up the online lib since some are not used at all.

GLT

  • The GLT deadtime contribution and low level corruption were discussed.

L3

  • Rainer and Arik are working on putting in the derivative change in the L3 tracking code. They also put in some conditions to make sure no loopers. The changes they made seems to improve the L3 tracking efficiency by about 4%. However, they are also confused by the fact that the number of reco tracks went down. They will add SVT hit variables to L1TNtuple to study this further.

Data Quality and Monitoring

  • Rahmat is working on OPR strip charts for the lumi ratio.

This page is created by Selina Li (selina@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.